A empty flask has a mass of 79.31 grams. A 25.0 mL sample of an unknown liquid was pipetted into the flask and gave a total mass (flask plus liquid) of 97.010 grams. The density of the unknown liquid was
1
Expert's answer
2018-09-28T06:17:43-0400
m(liquid) = total mass - mass of a empty flask = 97.010g - 79.31g = 17.7 g
density of the unknown liquid = m/V = 17.7g/25ml = 0.708 g/ml